# 111$\quad$1

### v1

@@ 3.1 Geometry Editing
On the geometry side, GaussianEditor [89] controls the 3DGS using the text prompts and semantic information from proposed Gaussian semantic tracing, which enables 3D inpainting, object removal, and composition for Gaussian Splatting, Gaussian Grouping [90] simultaneously rebuilds and segments open-world 3D objects under the supervision of 2D mask predictions from SAM and 3D spatial consistency constraints, which further enables diverse editing applications including 3D object removal, inpainting, and composition with high-quality visual effects and time efficiency. Furthermore, Point’n Move [91] combines interactive scene object manipulation with exposed region inpainting. Thanks to the explicit representation of 3DGS, the dual-stage self-prompting mask propagation process is proposed to transfer the given 2D prompt points to 3D mask segmentation, resulting in a user-friendly editing experience with high-quality effects. Although the above methods realize the editing on 3DGS, they are still limited to some simple editing operations (removal, rotation, and translation) for 3D objects. SuGaR [20] extracts explicit meshes from the 3DGS representation by regularizing Gaussians over surfaces. Further, it relies on manual adjustment of Gaussian parameters based on deformed meshes to realize desired geometry editing but struggles with large-scale deformation. SC-GS [26] learns a set of sparse control points for 3D scene dynamics but faces challenges with intense movements and detailed surface deformation. GaMeS [28] introduces a new GS-based model that combines conventional mesh and vanilla GS. The explicit mesh is utilized as input and parameterizes Gaussian components using the vertices, which can modify Gaussians in real time by altering mesh components during inference. However, it cannot handle significant deformations or changes, especially the deformation on large faces, since it cannot change the mesh topology during training. Although the above methods can finish some simple rigid transformations and non-rigid deformation, they still face challenges in their editing effectiveness and large-scale deformation. As shown in Fig. 7, Gao et al. [29] also adapt the mesh-based deformation to 3DGS by harnessing the priors of explicit representation (the surface properties like normals of the mesh, and the gradients generated by explicit deformation methods) and learning the face split to optimize the parameters and number of Gaussians, which provides adequate topological information to 3DGS and improves the quality for both the reconstruction and geometry editing results.
@@ 3.2 Appearance Editing
On the appearance side, GaussianEditor [92] proposes to first modify 2D images with language input with diffusion model [93] in the masked region generated by the recent 2D segmentation model [80] and updating attributes of Gaussians again similar to the previous NeRF editing work Instruct-NeRF2NeRF [94]. Another independent research work also named GaussianEditor [89] operates similarly but it further introduces a Hierarchical Gaussian Splatting (HGS) to allow 3D editing like object inpainting.
@@ 3.3 Physical Simulation
On the physical-based 3DGS editing, as shown in Fig. 9, PhysGaussian [19] employs discrete particle clouds from 3D GS for physically-based dynamics and photo-realistic rendering through continuum deformation [98] of Gaussian kernels. Gaussian Splashing [99] combines 3DGS and position-based dynamics (PBD) [100] to manage rendering, view synthesis, and solid/fluid dynamics cohesively. Similar to Gaussian shaders [22], the normal is applied to each Gaussian kernel to align its orientation with the surface normal and improve PBD simulation, also allowing the physically-based rendering to enhance dynamic surface reflections on fluids. VR-GS [27] is a physical dynamics-aware interactive Gaussian Splatting system for VR, tackling the difficulty of editing high-fidelity virtual content in real time. VR-GS utilizes 3DGS to close the quality gap between generated and manually crafted 3D content. By utilizing physically-based dynamics, which enhance immersion and offer precise interaction and manipulation controllability.
@@ 4.1 Segmentation and Understanding
NeRF-based methods are computationally intensive because of the implicit and continuous representation. Recent methods try to integrate 2D scene understanding methods with 3D Gaussians to produce a real-time and easy-to-editing 3D scene representation. Most methods utilize SAM [80] to produce semantic masks of input multi-view images [91, 90, 103, 104, 105, 106], or extract dense language features, CLIP [101]/DINO [102], of each pixel [107, 108, 109].
@@ 4.2 Geometry Reconstruction and SLAM
In the context of NeRF, a series of works [110, 111, 112, 113] have successfully reconstructed high-quality geometry from multi-view images. However, due to the discrete nature of 3DGS, only a few works stepped into this field. SuGaR [20] is the pioneering work that builds up 3D surfaces from multi-view images with the 3DGS representation. It introduces a simple but effective self-regularization loss to constrain that the distance between the camera and the closest Gaussian should be as close as possible to the corresponding pixel’s depth value in the rendered depth map, which encourages the alignment between 3DGS and the authentic 3D surface. Another work NeuSG [114] chooses to incorporate the previous NeRF-based surface reconstruction method NeuS [110] in the 3DGS representation to transfer the surface property to 3DGS. More specifically, it encourages Gaussians’ signed distance values to be zeros and the normal directions of 3DGS and the NeuS method to be as consistent as possible.
@@
There are also 3DGS methods targeting simultaneously localizing the cameras and reconstructing the 3D scenes. GS-SLAM [115] proposes an adaptive 3D Gaussian expanding strategy to add new 3D Gaussians into the training stage and delete unreliable ones with captured depths and rendered opacity values. To avoid duplicate densification, SplaTAM [116] uses view-independent colors for Gaussians and creates a densification mask to determine whether a pixel in a new frame needs densification by considering current Gaussians and the captured depth of the new frame. For stabilizing the localization and mapping, GaussianSplattingSLAM [117] and Gaussian-SLAM [118] put an extra scale regularization loss on the scale of Gaussians to encourage isotropic Gaussians. For easier initialization, LIV-GaussMap [119] initializes Gaussians with LiDAR point cloud and builds up an optimizable size-adaptive voxel grid for the global map. SGS-SLAM [120] further considers Gaussian’s semantic information in the simultaneous localization and mapping process by distilling 2D semantic information which can be obtained using 2D segmentation methods or provided by the dataset. We report quantitative results by different SLAM methods on the reconstruction task in Table. 4. The explicit geometry representation provided by 3DGS enables flexible reprojection to alleviate the misalignment between different viewpoints, thus leading to better reconstruction compared to NeRF-based methods.
@@ 4.3 Digital Human
For dynamic digital human modeling based on monocular video, GaussianBody [138] further leverages the physical-based priors to regularize the Guassians in the canonical space to avoid artifacts in the dynamic cloth from monocular video. GauHuman [139] re-designs the prune/split/clone of the original 3DGS to achieve efficient optimization and incorporates pose refinement and weight fields modules for fine details learning. It achieves minute-level training and real-time rendering (166 FPS). GaussianAvatar [140] incorporates the optimizable tensor with a dynamic appearance network to capture the dynamics better, allowing the dynamic avatar reconstruction and realistic novel animation in real time. Human101 [141] further pushes the speed of high-fidelity dynamic human creation to 100 seconds using a fixed-perspective camera.
@@
For human head modeling with 3DGS, MonoGaussianAvatar [136] first applies 3DGS to dynamic head reconstruction using the canonical space modeling and deformation prediction. Further, PSAvatar [142] introduces the explicit Flame face model [143] to initialize Gaussians, which can capture the high-fidelity facial geometry and even complicated volumetric objects (e.g.glasses). The tri-plane representation and the motion fields are used in GaussianHead [144] to simulate geometrically changing heads in continuous movements and render rich textures, including the skin and hair. For easier head expression controllability, GaussianAvatars [145] introduce the geometric priors (Flame parametric face model [143]) into 3DGS, which binds the Gaussians onto the explicit mesh and optimize the parameters of Gaussian ellipsoids. Rig3DGS [146] employs a learnable deformation to provide stability and generalization to novel expressions, head poses, and viewing directions to achieve controllable portraits on portable devices. In another way, HeadGas [147] attributes the 3DGS with a base of latent features that are weighted by the expression vector from 3DMMs [148], which achieves real-time animatable head reconstruction. FlashAvatar [149] further embeds a uniform 3D Gaussian field in a parametric face model and learns additional spatial offsets to capture facial details, successfully pushing the rendering speed to 300 FPS. To synthesize the high-resolution results, Gaussian Head Avatar [150] adopts the super-resolution network to achieve high-fidelity head avatar learning. Apart from these, some works extend the 3DGS into text-based head generation [151], DeepFake [152], and relighting [153].
@@
DreamGaussian [18] replaces the MipNeRF [38] representation in the DreamFusion [69] framework with 3DGS that uses SDS loss to optimize 3D Gaussians. The splitting process of 3DGS is very suitable for the optimization progress under the generative settings, so the efficiency advantages of 3DGS can be brought to text-to-3D generation based on the SDS loss. To improve the final quality, this work follows the idea of Magic3D [157] which extracts the mesh from the generated 3DGS and refines the texture details by optimizing UV textures through a pixel-wise Mean Squared Error (MSE) loss. In addition to 2D SDS, GSGEN [158] introduces a 3D SDS loss based on Point-E [159], a text-to-point-cloud diffusion model, to mitigate the multi-face or Janus problem. It adopts Point-E to initialize the point cloud as the initial geometry for optimization and also refines the appearance with only the 2D image prior. GaussianDreamer [160] also combines the priors of 2D and 3D diffusion models. It utilizes Shap-E [161] to generate the initial point cloud and optimizes 3DGS using 2D SDS. However, the generated initial point cloud is relatively sparse, so noisy point growth and color perturbation are further proposed to densify it. However, even if the 3D SDS loss is introduced, the Janus problem may still exist during optimization as the view is sampled one by one. Some methods [162, 163] fine-tune the 2D diffusion model [93] to generate multi-view images at once, thereby achieving multi-view supervision during SDS optimization. Or, the multi-view SDS proposed by BoostDream [164] directly creates a large $2\times 2$ images by stitching the rendered images from 4 sampled views and calculates the gradients under the condition of the multi-view normal map. This is a plug-and-play method that can first convert a 3D asset into differentiable representations including NeRF, 3DGS, and DMTet [165] through rendering supervision, and then optimize them to improve the quality of the 3D asset. Some methods have made improvements to SDS loss. LucidDreamer [166] proposes Interval Score Matching (ISM), which replaces DDPM in SDS with DDIM inversion and introduces supervisions from interval steps of the diffusion process to avoid high error in one-step reconstruction. Some generation results are shown in Fig. 12. GaussianDiffusion [167] proposes to incorporate structured noises from multiple viewpoints to alleviate the Janus problem and variational 3DGS for better generation results by mitigating floaters. Yang et al. [168] point out that the differences between the diffusion prior and the training process of the diffusion model will impair the quality of 3D generation, so they propose iterative optimization of the 3D model and the diffusion prior. Specifically, two additional learnable parameters are introduced in the classifier-free guidance formula, one is a learnable unconditional embedding, and the other is additional parameters added to the network, such as LoRA [169] parameters. These methods are not limited to 3DGS, and other originally NeRF-based methods including VSD [170], CSD [171] aiming at improving the SDS loss can be used for the 3DGS generation.
@@
The above methods focus on the generation of an individual object, while scene generation requires consideration of interactions and relationships between different objects. CG3D [179] inputs a text prompt manually deconstructed by the user into a scene graph, and the textual scene graph is interpreted as a probabilistic graphical model where the directed edge has a tail of the object node and a head of the interaction node. Then scene generation becomes an ancestor sampling by first generating the objects and then their interactions. The optimization is divided into two stages, where gravity and normal contact forces are introduced in the second stage. LucidDreamer [180] and Text2Immersion [181] are both based on a reference image (user-specified or text-generated) and extend outward to achieve 3D scene generation. The former utilizes Stable Diffusion (SD) [93] for image inpainting to generate unseen regions on the sampled views and incorporates monocular depth estimation and alignment to establish a 3D point cloud from these views. Finally, the point cloud is used as the initial value, and a 3DGS is trained using the projected images as the ground truth to achieve 3D scene generation. The latter method has a similar idea, while there is a process to remove outliers in the point cloud and the 3DGS optimization has two stages: training coarse 3DGS and refinement.
@@
Text-to-3D generation methods can be applied to image-to-3D generation, or monocular 3D reconstruction, with some simple modifications. For example, the pre-trained diffusion model used in SDS loss can be replaced with Zero-1-to-3 XL [182] for image condition [18]. We can also add losses between the input image and the corresponding rendered image under the input view to make the generation more consistent with the input image. Based on the image-to-3D generation of DreamGaussian [18], Repaint123 [183] proposes a progressive controllable repainting mechanism to refine the generated mesh texture. During the process of repainting the occlusions, it incorporates textural information from the reference image through attention feature injection [184] and proposes a visibility-aware repainting process to refine the overlap regions with different strengths. Finally, the refined images will be used as ground truths to directly optimize the texture through MSE loss, achieving fast optimization. Other methods explore utilizing existing 3D datasets [185, 186] and constructing large models to directly generate 3DGS representation from a single image. TriplaneGaussian [24] proposes a hybrid representation of tri-plane and 3DGS. It generates a point cloud and a tri-plane encoding 3DGS’s attributes information from the input image features through a transformer-based point cloud decoder and a tri-plane decoder, respectively. The generated point cloud is densified through an upsampling method and then projected onto the tri-plane to query features. The queried features are augmented by the projected image features and are translated into 3D Gaussian attributes using an MLP, thereby achieving the generation of 3DGS from a single image. AGG [187] also introduces a hybrid generator to obtain the point cloud and tri-plane features. But it first generates a coarse 3DGS and then upsamples it through a U-Net-based super-resolution module to improve the fidelity of generated results. LGM [188] first exploits the off-the-shelf models to generate multi-view images from text [162] or a single image [163]. Then it trains a U-Net-based network to generate 3DGS from the multi-view images. The U-Net is asymmetric, which allows for the input of high-resolution images while limiting the number of output Gaussians.

### v2

@@ 3.1 Geometry Editing
On the geometry side, GaussianEditor [134] controls the 3DGS using the text prompts and semantic information from proposed Gaussian semantic tracing, which enables 3D inpainting, object removal, and object composition. Gaussian Grouping [135] simultaneously rebuilds and segments open-world 3D objects under the supervision of 2D mask predictions from SAM and 3D spatial consistency constraints, which further enables diverse editing applications including 3D object removal, inpainting, and composition with high-quality visual effects and time efficiency. Furthermore, Point’n Move [136] combines interactive scene object manipulation with exposed region inpainting. Thanks to the explicit representation of 3DGS, the dual-stage self-prompting mask propagation process is proposed to transfer the given 2D prompt points to 3D mask segmentation, resulting in a user-friendly editing experience with high-quality effects. Feng et al. [137] propose a new Gaussian splitting algorithm to avoid inhomogeneous 3D Gaussian reconstruction and makes the boundary of 3D scenes after removal operation sharper. Although the above methods realize the editing on 3DGS, they are still limited to some simple editing operations (removal, rotation, and translation) for 3D objects. SuGaR [22] extracts explicit meshes from the 3DGS representation by regularizing Gaussians over surfaces. Further, it relies on manual adjustment of Gaussian parameters based on deformed meshes to realize desired geometry editing but struggles with large-scale deformation. SC-GS [28] learns a set of sparse control points for 3D scene dynamics but faces challenges with intense movements and detailed surface deformation. GaMeS [30] introduces a new GS-based model that combines conventional mesh and vanilla GS. The explicit mesh is utilized as input and parameterizes Gaussian components using the vertices, which can modify Gaussians in real time by altering mesh components during inference. However, it cannot handle significant deformations or changes, especially the deformation on large faces, since it cannot change the mesh topology during training. Although the above methods can finish some simple rigid transformations and non-rigid deformation, they still face challenges in their editing effectiveness and large-scale deformation. As shown in Fig. 7, Gao et al. [31] also adapt the mesh-based deformation to 3DGS by harnessing the priors of explicit representation (the surface properties like normals of the mesh, and the gradients generated by explicit deformation methods) and learning the face split to optimize the parameters and number of Gaussians, which provides adequate topological information to 3DGS and improves the quality for both the reconstruction and geometry editing results. GaussianFrosting [138] shares a similar idea as Gao et al. [31] by constructing a base mesh but further develops a forsting layer to allow Gaussians to move in a small range near the mesh surface.
@@ 3.2 Appearance Editing
On the appearance side, GaussianEditor [139] proposes to first modify 2D images with language input with diffusion model [140] in the masked region generated by the recent 2D segmentation model [110] and updating attributes of Gaussians again similar to the previous NeRF editing work Instruct-NeRF2NeRF [141]. Another independent research work also named GaussianEditor [134] operates similarly but it further introduces a Hierarchical Gaussian Splatting (HGS) to allow 3D editing like object inpainting. GSEdit [142] takes a texture mesh or pre-trained 3DGS as input and utilizes the Instruct-Pix2Pix [143] and the SDS loss to updated the input mesh or 3DGS. To alleviate the inconsistency issue, GaussCtrl [144] introduces the depth map as the conditional input of the ControlNet [99] to encourage geometry consistency. Wang et al. [145] also aims to solve this inconsistent issue by introducing multi-view cross-attention maps. Texture-GS [146] proposes to disentangle the geometry and appearance of 3DGS and learns a UV mapping network for points near the underlying surface, thus enabling manipulations including texture painting and texture swapping. 3DGM [147] also represents a 3D scene with a proxy mesh with fixed UV mapping where Gaussians are stored on the texture map. This disentangled representation also allows animation and texture editing. Apart from local texture editing, there are works [148, 149, 150] focusing on stylizing 3DGS with a reference style image.
@@ 3.3 Physical Simulation
On the physical-based 3DGS editing, as shown in Fig. 9, PhysGaussian [21] employs discrete particle clouds from 3D GS for physically-based dynamics and photo-realistic rendering through continuum deformation [154] of Gaussian kernels. Gaussian Splashing [155] combines 3DGS and position-based dynamics (PBD) [156] to manage rendering, view synthesis, and solid/fluid dynamics cohesively. Similar to Gaussian shaders [24], the normal is applied to each Gaussian kernel to align its orientation with the surface normal and improve PBD simulation, also allowing the physically-based rendering to enhance dynamic surface reflections on fluids. VR-GS [29] is a physical dynamics-aware interactive Gaussian Splatting system for VR, tackling the difficulty of editing high-fidelity virtual content in real time. VR-GS utilizes 3DGS to close the quality gap between generated and manually crafted 3D content. By utilizing physically-based dynamics, which enhance immersion and offer precise interaction and manipulation controllability. Spring-Gaus [157] applies the Spring-Mass model into the modeling of dynamic 3DGS and learns the physical properties like mass and velocity from input video, which are editable for real-world simulation. Feature Splatting [158] further incorporates semantic priors from pre-trained networks and makes object-level simulation possible.
@@ 4.1 Segmentation and Understanding
NeRF-based methods are computationally intensive because of the implicit and continuous representation. Recent methods try to integrate 2D scene understanding methods with 3D Gaussians to produce a real-time and easy-to-editing 3D scene representation. Most methods utilize pre-trained 2D segmentation methods like SAM [110] to produce semantic masks of input multi-view images [136, 135, 161, 162, 163, 164, 165, 166], or extract dense language features, CLIP [159]/DINO [160], of each pixel [167, 168, 169].
@@ 4.2 Geometry Reconstruction and SLAM
In the context of NeRF, a series of works [170, 171, 172, 173] have successfully reconstructed high-quality geometry from multi-view images. However, due to the discrete nature of 3DGS, only a few works stepped into this field. SuGaR [22] is the pioneering work that builds up 3D surfaces from multi-view images with the 3DGS representation. It introduces a simple but effective self-regularization loss to constrain that the distance between the camera and the closest Gaussian should be as close as possible to the corresponding pixel’s depth value in the rendered depth map, which encourages the alignment between 3DGS and the authentic 3D surface. Another work NeuSG [174] chooses to incorporate the previous NeRF-based surface reconstruction method NeuS [170] in the 3DGS representation to transfer the surface property to 3DGS. More specifically, it encourages Gaussians’ signed distance values to be zeros and the normal directions of 3DGS and the NeuS method to be as consistent as possible. 3DGSR [175] and GSDF [176] also encourage the consistency between SDF and 3DGS to enhance the geometry reconstruction quality. DN-Splatter [177] utilizes depth and normal priors captured from common devices or predicted from general-purpose networks to enhance the reconstruction quality of 3DGS. Wolf et al. [178] first train a 3DGS to render stereo-calibrated novel views and apply stereo depth estimation on the rendered views. The estimated dense depth maps are fused together by the Truncated Signed Distance Function(TSDF) to form a triangular mesh. 2D-GS [179] replaces 3D Gaussians with 2D Gaussians for more accurate ray-splat intersection and employs a low-pass filter to avoid degenerated line projection. Though attempts have been made in the field of geometry reconstruction, due to the discrete nature of 3DGS, current methods achieve comparable or worse results compared to implicit representation-based methods with continuous field assumption where the surface can be easily determined.
@@
There are also 3DGS methods targeting simultaneously localizing the cameras and reconstructing the 3D scenes. GS-SLAM [180] proposes an adaptive 3D Gaussian expanding strategy to add new 3D Gaussians into the training stage and delete unreliable ones with captured depths and rendered opacity values. To avoid duplicate densification, SplaTAM [181] uses view-independent colors for Gaussians and creates a densification mask to determine whether a pixel in a new frame needs densification by considering current Gaussians and the captured depth of the new frame. For stabilizing the localization and mapping, GaussianSplattingSLAM [182] and Gaussian-SLAM [183] put an extra scale regularization loss on the scale of Gaussians to encourage isotropic Gaussians. For easier initialization, LIV-GaussMap [184] initializes Gaussians with LiDAR point cloud and builds up an optimizable size-adaptive voxel grid for the global map. SGS-SLAM [185], NEDS-SLAM [186], and SemGauss-SLAM [187] further consider Gaussian’s semantic information in the simultaneous localization and mapping process by distilling 2D semantic information which can be obtained using 2D segmentation methods or provided by the dataset. Deng et al. [188] avoid redundant Gaussian splitting based on sliding window mask and use vector quantization to further encourage compact 3DGS. CG-SLAM [189] introduces an uncertainty map into the training process based on the rendered depth and greatly improves the reconstruction quality. Based on the reconstructed map by SLAM-based methods, tasks in robotics like relocalization [190], navigation [191, 192, 193], 6D pose estimation [194], multi-sensor calibration [195, 196] and manipulation [197, 198] can be performed efficiently. We report quantitative results by different SLAM methods on the reconstruction task in Table. 4. The explicit geometry representation provided by 3DGS enables flexible reprojection to alleviate the misalignment between different viewpoints, thus leading to better reconstruction compared to NeRF-based methods. The real-time rendering feature of 3DGS also makes neural-based SLAM methods more applicable as the training in previous NeRF-based methods require more hardware and time.
@@ 4.3 Digital Human
For dynamic digital human modeling based on monocular video, GaussianBody [216] further leverages the physical-based priors to regularize the Guassians in the canonical space to avoid artifacts in the dynamic cloth from monocular video. GauHuman [217] re-designs the prune/split/clone of the original 3DGS to achieve efficient optimization and incorporates pose refinement and weight fields modules for fine details learning. It achieves minute-level training and real-time rendering (166 FPS). GaussianAvatar [218] incorporates the optimizable tensor with a dynamic appearance network to capture the dynamics better, allowing the dynamic avatar reconstruction and realistic novel animation in real time. Human101 [219] further pushes the speed of high-fidelity dynamic human creation to 100 seconds using a fixed-perspective camera. Simliar to [31], SplattingAvatar [220] and GoMAvatar [221] embed Gaussians onto a canonical human body mesh. The positions of a Gaussian is determined by the barycentric point and the displacement along the normal direction. To resolve the unbalanced aggregation of Gaussians caused by densification and splitting operation, GVA [222] proposes a surface-guided Gaussians re-initialization strategy to make the trained Gaussians better fit the input monocular video. HAHA [223] also attaches Gaussians onto the surface of a mesh but combines rendered results from a textured human body mesh and Gaussians together to reduce the number of Gaussians.
@@
For human head modeling with 3DGS, MonoGaussianAvatar [214] first applies 3DGS to dynamic head reconstruction using the canonical space modeling and deformation prediction. Further, PSAvatar [224] introduces the explicit Flame face model [225] to initialize Gaussians, which can capture the high-fidelity facial geometry and even complicated volumetric objects (e.g.glasses). The tri-plane representation and the motion fields are used in GaussianHead [226] to simulate geometrically changing heads in continuous movements and render rich textures, including the skin and hair. For easier head expression controllability, GaussianAvatars [227] introduce the geometric priors (Flame parametric face model [225]) into 3DGS, which binds the Gaussians onto the explicit mesh and optimize the parameters of Gaussian ellipsoids. Rig3DGS [228] employs a learnable deformation to provide stability and generalization to novel expressions, head poses, and viewing directions to achieve controllable portraits on portable devices. In another way, HeadGas [229] attributes the 3DGS with a base of latent features that are weighted by the expression vector from 3DMMs [230], which achieves real-time animatable head reconstruction. FlashAvatar [231] further embeds a uniform 3D Gaussian field in a parametric face model and learns additional spatial offsets to capture facial details, successfully pushing the rendering speed to 300 FPS. To synthesize the high-resolution results, Gaussian Head Avatar [232] adopts the super-resolution network to achieve high-fidelity head avatar learning. To synthesize high-quality avatar from few-view input, SplatFace [233] propose to first initialize Gaussians on a template mesh and jointly optimizing Gaussians and the mesh with a splat-to-mesh distance loss. GauMesh [234] propose a hybrid representation containing both tracked textured meshes and canonical 3D Gaussians together with a learnable deformation field to represent dynamic human head. Apart from these, some works extend the 3DGS into text-based head generation [235], DeepFake [236], and relighting [237].
@@
DreamGaussian [20] replaces the MipNeRF [52] representation in the DreamFusion [97] framework with 3DGS that uses SDS loss to optimize 3D Gaussians. The splitting process of 3DGS is very suitable for the optimization progress under the generative settings, so the efficiency advantages of 3DGS can be brought to text-to-3D generation based on the SDS loss. To improve the final quality, this work follows the idea of Magic3D [241] which extracts the mesh from the generated 3DGS and refines the texture details by optimizing UV textures through a pixel-wise Mean Squared Error (MSE) loss. In addition to 2D SDS, GSGEN [242] introduces a 3D SDS loss based on Point-E [243], a text-to-point-cloud diffusion model, to mitigate the multi-face or Janus problem. It adopts Point-E to initialize the point cloud as the initial geometry for optimization and also refines the appearance with only the 2D image prior. GaussianDreamer [244] also combines the priors of 2D and 3D diffusion models. It utilizes Shap-E [245] to generate the initial point cloud and optimizes 3DGS using 2D SDS. However, the generated initial point cloud is relatively sparse, so noisy point growth and color perturbation are further proposed to densify it. However, even if the 3D SDS loss is introduced, the Janus problem may still exist during optimization as the view is sampled one by one. Some methods [246, 247] fine-tune the 2D diffusion model [140] to generate multi-view images at once, thereby achieving multi-view supervision during SDS optimization. Or, the multi-view SDS proposed by BoostDream [248] directly creates a large $2\times 2$ images by stitching the rendered images from 4 sampled views and calculates the gradients under the condition of the multi-view normal map. This is a plug-and-play method that can first convert a 3D asset into differentiable representations including NeRF, 3DGS, and DMTet [249] through rendering supervision, and then optimize them to improve the quality of the 3D asset. Some methods have made improvements to SDS loss. LucidDreamer [250] proposes Interval Score Matching (ISM), which replaces DDPM in SDS with DDIM inversion and introduces supervisions from interval steps of the diffusion process to avoid high error in one-step reconstruction. Some generation results are shown in Fig. 12. GaussianDiffusion [251] proposes to incorporate structured noises from multiple viewpoints to alleviate the Janus problem and variational 3DGS for better generation results by mitigating floaters. Yang et al. [252] point out that the differences between the diffusion prior and the training process of the diffusion model will impair the quality of 3D generation, so they propose iterative optimization of the 3D model and the diffusion prior. Specifically, two additional learnable parameters are introduced in the classifier-free guidance formula, one is a learnable unconditional embedding, and the other is additional parameters added to the network, such as LoRA [253] parameters. These methods are not limited to 3DGS, and other originally NeRF-based methods including VSD [254] and CSD [255] aiming at improving the SDS loss can be used for the 3DGS generation. GaussianCube [256] instead trains a 3D diffusion model based on a GaussianCube representation that is converted from a constant number of Gaussians with voxelization via optimal transport. GVGEN [257] also works in 3D space but is based on a 3D Gaussian volume representation.
@@
The above methods focus on the generation of an individual object, while scene generation requires consideration of interactions and relationships between different objects. CG3D [265] inputs a text prompt manually deconstructed by the user into a scene graph, and the textual scene graph is interpreted as a probabilistic graphical model where the directed edge has a tail of the object node and a head of the interaction node. Then scene generation becomes an ancestor sampling by first generating the objects and then their interactions. The optimization is divided into two stages, where gravity and normal contact forces are introduced in the second stage. LucidDreamer [266] and Text2Immersion [267] are both based on a reference image (user-specified or text-generated) and extend outward to achieve 3D scene generation. The former utilizes Stable Diffusion (SD) [140] for image inpainting to generate unseen regions on the sampled views and incorporates monocular depth estimation and alignment to establish a 3D point cloud from these views. Finally, the point cloud is used as the initial value, and a 3DGS is trained using the projected images as the ground truth to achieve 3D scene generation. The latter method has a similar idea, while there is a process to remove outliers in the point cloud and the 3DGS optimization has two stages: training coarse 3DGS and refinement. To explore 3D scene generation, GALA3D [268] leverages both the object-level text-to-3D model MVDream [246] to generate realistic objects and scene-level diffusion model to compose them. DreamScene [269] proposes multi-timestep sampling for multi-stage scene-level generation to synthesize surrounding environment, ground, and objects to avoid complicate object composition. Instead of separately modeling each object, RealmDreamer [270] utilizes the diffusion priors from inpainting and depth estimation models to generate different viewpoints of a scene iteratively. DreamScene360 [271] instead generates a 360-degree panoramic image and converts it into a 3D scene with depth estimation.
@@
Text-to-3D generation methods can be applied to image-to-3D generation, or monocular 3D reconstruction, with some simple modifications. For example, the pre-trained diffusion model used in SDS loss can be replaced with Zero-1-to-3 XL [272] for image condition [20]. We can also add losses between the input image and the corresponding rendered image under the input view to make the generation more consistent with the input image. Based on the image-to-3D generation of DreamGaussian [20], Repaint123 [273] proposes a progressive controllable repainting mechanism to refine the generated mesh texture. During the process of repainting the occlusions, it incorporates textural information from the reference image through attention feature injection [274] and proposes a visibility-aware repainting process to refine the overlap regions with different strengths. Finally, the refined images will be used as ground truths to directly optimize the texture through MSE loss, achieving fast optimization. Other methods explore utilizing existing 3D datasets [275, 276] and constructing large models to directly generate 3DGS representation from a single image. TriplaneGaussian [26] proposes a hybrid representation of tri-plane and 3DGS. It generates a point cloud and a tri-plane encoding 3DGS’s attributes information from the input image features through a transformer-based point cloud decoder and a tri-plane decoder, respectively. The generated point cloud is densified through an upsampling method and then projected onto the tri-plane to query features. The queried features are augmented by the projected image features and are translated into 3D Gaussian attributes using an MLP, thereby achieving the generation of 3DGS from a single image. LGM [277] first exploits the off-the-shelf models to generate multi-view images from text [246] or a single image [247]. Then it trains a U-Net-based network to generate 3DGS from the multi-view images. The U-Net is asymmetric, which allows for the input of high-resolution images while limiting the number of output Gaussians. AGG [278] also introduces a hybrid generator to obtain the point cloud and tri-plane features. But it first generates a coarse 3DGS and then upsamples it through a U-Net-based super-resolution module to improve the fidelity of generated results. Instead of predicting point cloud from image, BrightDreamer [279] predicts the deviation of a set of fixed anchor points to determine the centers of Gaussians. GRM [280] utilizes existing multi-view generation models to train a pixel-aligned Gaussian representation for faithful 3D generation via a single feed-forward pass. IM-3D [281] finetunes the image-to-video model based on Emu [282] to generate a turn-table like video rotating around an object, which are taken as the input for 3DGS reconstruction. Gamba [283] proposes to predict Gaussian attributes with the recent Mamba [284] network to better capture the relationship between Gaussians. MVControl [285] extends ControlNet [99] to 3D generation task and allows extra condition inputs like edge, depth, normal, and scribble fed into existing multi-view generation models. Hyper-3DG [286] proposes a geometry and texture refinement module to improve generation quality based on hypergraph learning where each node is a patch image of a coarse 3DGS. For the same purpose, DreamPolisher [287] utilizes a ControlNet-based network for texture refinement and ensures consistency between different viewpoints with a view-consistent geometric guidance. FDGaussian [288] injects tri-plane features into the diffusion model for better geometry generation.